# UniMS-RAG: A Unified Multi-source Retrieval-Augmented Generation for Personalized Dialogue Systems

### v2

@@ 1 Introduction
2 Related Work
2.1 Personalized Dialogue System
@@
2.2 Knowledge-grounded Dialogue System
@@
2.3 Retrieval-augmented Generation
@@ 4.1.2 Retrieval
To joint train the retriever and reader in a unified manner, we first reformulate the traditional classification task (relevant or irrelevant) for retrieval into a generation task. Then we define several similarity tokens as shown in Table I, ranging from 0 to 1 with the internal as 0.1. In this way, we enforce the UniMS-RAG to predict the similarity score by generating corresponding tokens. To get the supervised label of similarity scores, there are two different ways: 1) the similarity scores are generated by independent trained retriever such as DPR [42]; 2) the similarity scores are generated by prompting LLMs [45]. The details can be found in § 4.2. Furthermore, we design a evidence attention mask mechanism to mask unrelated evidence when predicting the similarity score for current evidence, aiming to reduce unnecessary noises in the context as shown in Figure 2.
@@ 4.1.3 Generation
. We use [EVIDENCE] and [EOE] to represent the start and end positions of the retrieved evidences. The $[Sim_{i}]$ stands for the similarity score of $i_{th}$ evidence, calculated using the retriever (§4.2).
4.2 General Framework
@@
4.2.1 Relevance Scores Acquisition
@@ 4.2.2 Training Stage
4.2.3 Inference Stage
In this section, we introduce how to refine the generated responses, considering the similarity score between dialogue context and retrieved evidences and consistency score between retrieved evidences and generated responses. Algorithm 1 shows the details of the whole processing. Specifically, we first calculate the consistency score between each retrieved evidence ($e_{i}$) and current generated response ($r$) using a fine-tuned natural language inference model (see § 5.4). Next, we determine the overall score for each evidence by multiplying its similarity score with the consistency score. This combined score serves as an indicator of the evidence’s quality, taking into account its relevance to the dialogue context and its alignment with the generated response. In this way, we can adjust the evidences based on their determined quality and a predefined update number, resulting in the creation of a new list of evidences. This process ensures that the most relevant and high-quality pieces of evidence are given priority, contributing to the refinement of the overall generated responses. Ultimately, we can re-generate the responses $r_{new}$ using a new list of evidences, providing the dialogue context. Moreover, we can iteratively apply the self-refinement algorithm by incorporating the newly generated responses, denoted as $r_{new}$, and the updated list of evidences. We provide a thorough analysis at § 7.2, outlining the rationale behind the selection of update numbers and the procedural steps involved in refining the responses.
@@ 4.2.3 Inference Stage
5 Experimental Setups
5.1 Research Questions
@@ 5.2 Datasets
5.3 Baselines and Evaluation Metrics
5.3.1 Knowledge Source Selection
@@ 5.3.3 Response Generation
5.4 Implementation Details
UniMS-RAG: We mainly choose ChatGLM-6B [55, 56] as the backbone models during training, we set the batch size as 8, train models with 3 epochs and save the checkpoint with the lowest validation loss. For other hyper-parameter settings, we mainly follow the corresponding official code  [^4] https://github.com/THUDM/ChatGLM-6B. Due to the computation limit, we conduct training with LoRA [57] at one single 3090 GPU, and it costs about 4-6 hours. For the prompting using LLMs, we choose gpt-3.5-turbo-1106 and gpt-4o, and set both the temperature and top p as 0.1 to reduce the randomness of LLMs. We choose only the top-ranked result for the main experiment to ensure a consistent setting, and we additionally investigate the effects of retrieving different numbers of results to evaluate whether UniMS-RAG can effectively filter out unrelated evidence. According to different sources of similarity scores, we have three variants: UniMS-RAG (ChatGPT), UniMS-RAG (GPT-4o) and UniMS-RAG (DPR) where the similarity scores come from ChatGPT, GPT4o and DPR, respectively. At the inference stage, we use self-predicted similarity scores to retrieve the evidence, serving as an indicator. More variants or analyses can be found in §7.2.
@@ 6.2.2 The performance of UniMS-RAG
TABLE V: The performance of Generation of different methods, in which part of the results are directly copied from wang2023large [11]. We follow the official code of SAFARI to conduct evaluation on DuLeMon dataset. We bold the highest performance and underline the second-best performance.
@@ 6.3.2 The performance of UniMS-RAG
6.3.3 Discussion of Unified Manner
@@
7 Analysis and Discussions
In this section, we choose the best model as shown in our previous experiments to investigate the performance changes under different settings (UniMS-RAG w/ DPR). In detail, we start from the performance of our proposed model with different numbers of retrieved evidence (§7.1), then we study the effects of introduced self-refinement mechanism during the inference stage by re-evaluating the relationship between generated response with the dialogue context or retrieved evidence (§7.2). We present the ablation study to show the effectiveness and rationale of UniMS-RAG (§7.2), followed by the results of human evaluation and case study (See supplementary).
7.1 Different Numbers of Retrieved Results
@@
7.2 Additional Self-refinement during Inference
@@ 7.3 Ablation Study
7.4 Human Evaluation
Human evaluation is conducted to evaluate the quality of generated response in terms of three metrics: coherence score (Coh.), persona consistency score (Per.Cs for both DuLeMon and KBP), and knowledge consistency score (Know.Cs only for KBP). We randomly sample 100 responses from each dataset with grounding information for each model (We choose UniMS-RAG w/ DPR or w/ ChatGPT and using itself as retriever) and ask three well-educated annotators [^6] The human inspection and annotation were conducted by a reputable data annotation company, and the annotators are compensated fairly based on the market price without revealing any personal information to indicate its coherence score (1-5) and whether the response is consistent with the given persona (1/0), and knowledge (1/0). Table VIII shows the final result. When the number of retrieved evidence is one, we observe there is a significant improvement on DuLeMon and a slight improvement on KBP datasets when using UniMS-RAG w/ DPR. Moreover, when the number of retrieved evidence increases, it is obvious that the performance is largely improved. This observation reveals the effectiveness and robustness of UniMS-RAG to denoise unrelated information, and capture relevant evidences, thanks to the introduction of task and unique training strategies. In addition, UniMS-RAG w/ ChatGPT performs slightly worse than UniMS-RAG w/ DPR due to worse similarity signals. We also observe that human is more likely to find persona-inconsistent cases [11]. There are some responses that have intra-sentence or intra-context contradictions [58], for example, the dialogue responds: ”Yes, our license plate starts with ‘Yun A’” to the last user turn: ”Does your license plate start with ‘Yun B’?”.
@@ 7.4 Human Evaluation
7.5 Case Study
Figure 6: The case study (No cherry-picking). We use orange to indicate acting tokens and blue to indicate evaluation tokens.
To better understand how the proposed UniMS-RAG works, Figure 6 presents a specific dialogue from KBP. At the current turn, the user inquires about the national language of the personalized dialogue system, and the UniMS-RAG determine to call PERSONA source first and then the corresponding DOCUMENT to answer the inquiry. This is correct according to the ground-truth labels in the original dataset. Then using itself as retriever to retrieve top-3 evidences from each source sequentially. However, in this task, we can find that 1) the highly related persona descriptions only get 0.5 in terms of relevance despite it still is the highest; 2) there are several same relevance score of evidences in DOCUMENT source since it is relatively difficult for the model to distinguish the differences. We consider this is also related to the hard negative mining during the retriever training. Finally, given the context, source signals and corresponding evidences and relevance scores, UniMS-RAG is capable to generate persona- and knowledge-consistent responses. In this way, UniMS-RAG can serve as planner, retriever, and reader in a unified manner.
7.6 Error Analysis and Limitations
Despite the effectiveness and robustness of the proposed UniMS-RAG framework for PerDS, we would like to better understand the bottleneck and limitations of UniMS-RAG for further improvement in future studies. After analyzing those fail cases, we identify the following limitations and discuss the potential directions:
  • Error Propagation. This is a typical issue of solving multiple tasks in sequential order. As shown in Table VII, there is still a gap (i.e., 15% - 20%) between the original UniMS-RAG and the one providing ground-truth planning and retrieval. To avoid this issue, there are some techniques studied currently, including contrastive objectives [59] and noisy channel models [60].
  • Better Planning and Retrieval. As shown in Table V and analysis, better planning and retrieval always lead to better performance, such as the increased number of retrieved evidences, and additional self-refinement mechanism. We also successfully demonstrate that LLMs themselves can be used as retriever simultaneously. Since we observe UniMS-RAG performs dramatically with different dialogue data and different similarity signals, several possible directions are to study how to get more large scale high-quality data or investigate different interactions (i.e, more complex prompting) [61, 35] to further enhance the performance of planning and retrieval.
  • The Number and Organization of Different Sources. We mainly target on four sources of knowledge in this paper (i.e, User-Per, Bot-Per, DOCUMENTS, NULL), however, there are many other useful sources in practice [30], such as user memory [62], local database [63] and so on. In addition, the organization of different sources of knowledge also plays a key role in the RAG, such as the different data format [25, 64], or different relationships within sources [35].
8 Conclusion
@@
Acknowledgments
This paper was partially supported by grants from the RGC General Research Funding Scheme (GRF) 14222922 (CUHK 2151185). Work done when Hongru Wang is visiting The University of Edinburgh.

### v3

@@ I Introduction
II Related Work
II-A Personalized Dialogue System
@@
II-B Knowledge-grounded Dialogue System
@@
II-C Retrieval-augmented Generation
@@ IV-A2 Retrieval
To joint train the retriever and reader in a unified manner, we first reformulate the traditional classification task (relevant or irrelevant) for retrieval into a generation task. Then we define several similarity tokens as shown in Table I, ranging from 0 to 1 with the internal as 0.1. In this way, we enforce the UniMS-RAG to predict the similarity score by generating corresponding tokens. To get the supervised label of similarity scores, there are two different ways: 1) the similarity scores are generated by independent trained retriever such as DPR [42]; 2) the similarity scores are generated by prompting LLMs [45]. The details can be found in § IV-B. Furthermore, we design a evidence attention mask mechanism to mask unrelated evidence when predicting the similarity score for current evidence, aiming to reduce unnecessary noises in the context as shown in Figure 2.
@@ IV-A3 Generation
. We use [EVIDENCE] and [EOE] to represent the start and end positions of the retrieved evidences. The $[Sim_{i}]$ stands for the similarity score of $i_{th}$ evidence, calculated using the retriever (§IV-B).
IV-B General Framework
@@
IV-B1 Relevance Scores Acquisition
@@ IV-B2 Training Stage
IV-B3 Inference Stage
In this section, we introduce how to refine the generated responses, considering the similarity score between dialogue context and retrieved evidences and consistency score between retrieved evidences and generated responses. Algorithm 1 shows the details of the whole processing. Specifically, we first calculate the consistency score between each retrieved evidence ($e_{i}$) and current generated response ($r$) using a fine-tuned natural language inference model (see § V-D). Next, we determine the overall score for each evidence by multiplying its similarity score with the consistency score. This combined score serves as an indicator of the evidence’s quality, taking into account its relevance to the dialogue context and its alignment with the generated response. In this way, we can adjust the evidences based on their determined quality and a predefined update number, resulting in the creation of a new list of evidences. This process ensures that the most relevant and high-quality pieces of evidence are given priority, contributing to the refinement of the overall generated responses. Ultimately, we can re-generate the responses $r_{new}$ using a new list of evidences, providing the dialogue context. Moreover, we can iteratively apply the self-refinement algorithm by incorporating the newly generated responses, denoted as $r_{new}$, and the updated list of evidences. We provide a thorough analysis at § VII-B, outlining the rationale behind the selection of update numbers and the procedural steps involved in refining the responses.
@@ IV-B3 Inference Stage
V Experimental Setups
V-A Research Questions
@@ V-B Datasets
V-C Baselines and Evaluation Metrics
V-C1 Knowledge Source Selection
@@ V-C3 Response Generation
V-D Implementation Details
UniMS-RAG: We mainly choose ChatGLM-6B [55, 56] as the backbone models during training, we set the batch size as 8, train models with 3 epochs and save the checkpoint with the lowest validation loss. For other hyper-parameter settings, we mainly follow the corresponding official code  [^4] https://github.com/THUDM/ChatGLM-6B. Due to the computation limit, we conduct training with LoRA [57] at one single 3090 GPU, and it costs about 4-6 hours. For the prompting using LLMs, we choose gpt-3.5-turbo-1106 and gpt-4o, and set both the temperature and top p as 0.1 to reduce the randomness of LLMs. We choose only the top-ranked result for the main experiment to ensure a consistent setting, and we additionally investigate the effects of retrieving different numbers of results to evaluate whether UniMS-RAG can effectively filter out unrelated evidence. According to different sources of similarity scores, we have three variants: UniMS-RAG (ChatGPT), UniMS-RAG (GPT-4o) and UniMS-RAG (DPR) where the similarity scores come from ChatGPT, GPT4o and DPR, respectively. At the inference stage, we use self-predicted similarity scores to retrieve the evidence, serving as an indicator. More variants or analyses can be found in §VII-B.
@@ VI-B2 The performance of UniMS-RAG
TABLE V: The performance of Generation of different methods. We follow the official code of SAFARI to conduct evaluation on DuLeMon dataset. We bold the highest performance and underline the second-best performance.
@@ VI-C2 The performance of UniMS-RAG
VI-C3 Discussion of Unified Manner
@@
VII Analysis and Discussions
In this section, we choose the best model as shown in our previous experiments to investigate the performance changes under different settings (UniMS-RAG w/ DPR). In detail, we start from the performance of our proposed model with different numbers of retrieved evidence (§VII-A), then we study the effects of introduced self-refinement mechanism during the inference stage by re-evaluating the relationship between generated response with the dialogue context or retrieved evidence (§VII-B). We present the ablation study to show the effectiveness and rationale of UniMS-RAG (§VII-B), followed by the results of human evaluation.
VII-A Different Numbers of Retrieved Results
@@
VII-B Additional Self-refinement during Inference
@@ VII-C Ablation Study
VII-D Human Evaluation
Human evaluation is conducted to evaluate the quality of generated response in terms of three metrics: coherence score (Coh.), persona consistency score (Per.Cs for both DuLeMon and KBP), and knowledge consistency score (Know.Cs only for KBP). We randomly sample 100 responses from each dataset with grounding information for each model (We choose UniMS-RAG w/ DPR or w/ ChatGPT and using itself as retriever) and ask three well-educated annotators [^6] The human inspection and annotation were conducted by a reputable data annotation company, and the annotators are compensated fairly based on the market price without revealing any personal information to indicate its coherence score (1-5) and whether the response is consistent with the given persona (1/0), and knowledge (1/0). Table VIII shows the final result. When the number of retrieved evidence is one, we observe there is a significant improvement on DuLeMon and a slight improvement on KBP datasets when using UniMS-RAG w/ DPR. Moreover, when the number of retrieved evidence increases, it is obvious that the performance is largely improved. This observation reveals the effectiveness and robustness of UniMS-RAG to denoise unrelated information, and capture relevant evidences, thanks to the introduction of task and unique training strategies. In addition, UniMS-RAG w/ ChatGPT performs slightly worse than UniMS-RAG w/ DPR due to worse similarity signals. We also observe that human is more likely to find persona-inconsistent cases [11]. There are some responses that have intra-sentence or intra-context contradictions [58], for example, the dialogue responds: “Yes, our license plate starts with ‘Yun A’” to the last user turn: “Does your license plate start with ‘Yun B’?”.
@@ VII-D Human Evaluation
VIII Conclusion
@@
Acknowledgment
This paper was partially supported by grants from the RGC General Research Funding Scheme (GRF) 14222922 (CUHK 2151185). Work done when Hongru Wang is visiting EdinburghNLP.